### v1

# Adversarial Training for Multimodal Large Language Models against Jailbreak Attacks

Abstract
Multimodal large language models (MLLMs) have made remarkable strides in cross-modal comprehension and generation tasks. However, they remain vulnerable to jailbreak attacks, where crafted perturbations bypass security guardrails and elicit harmful outputs. In this paper, we present the first adversarial training (AT) paradigm tailored to defend against jailbreak attacks during the MLLM training phase. Extending traditional AT to this domain poses two critical challenges: efficiently tuning massive parameters and ensuring robustness against attacks across multiple modalities. To address these challenges, we introduce Projection Layer Against Adversarial Training (ProEAT), an end-to-end AT framework. ProEAT incorporates a projector-based adversarial training architecture that efficiently handles large-scale parameters while maintaining computational feasibility by focusing adversarial training on a lightweight projector layer instead of the entire model; additionally, we design a dynamic weight adjustment mechanism that optimizes the loss function’s weight allocation based on task demands, streamlining the tuning process. To enhance defense performance, we propose a joint optimization strategy across visual and textual modalities, ensuring robust resistance to jailbreak attacks originating from either modality. Extensive experiments conducted on five major jailbreak attack methods across three mainstream MLLMs demonstrate the effectiveness of our approach. ProEAT achieves state-of-the-art defense performance, outperforming existing baselines by an average margin of +34% across text and image modalities, while incurring only a 1% reduction in clean accuracy. Furthermore, evaluations on real-world embodied intelligent systems highlight the practical applicability of our framework, paving the way for the development of more secure and reliable multimodal systems.
Figure 1: Overview of our joint multimodal optimization framework. The added optimization enables MLLMs to improve their defense capabilities, ultimately resulting in reduced vulnerability to jailbreak attacks.
1 Introduction
Equal contribution.
Multimodal large language models (MLLMs) [2, 3, 17, 21, 30] have achieved remarkable success in various task, including text-to-image generation [59, 12], visual question answering [26, 24], and multi-turn dialogues [14, 51]. Notable examples, such as GPT-4 [1] and LLaVA [43], have demonstrated exceptional capabilities, particularly after being fine-tuned with instructions and aligned with human feedback. These advancements are fueled by innovations in data scale, computational resources, and model architectures. By extending the functionality of Large Language Models (LLMs) [33, 9, 42], MLLMs have broadened their utility in diverse domains.
However, a critical concern is that this integration of multimodal capabilities introduces significant new security challenges, particularly susceptibility to cross-modal jailbreak attacks [28, 47, 40, 63]. Attackers maliciously craft multimodal inputs for inappropriate content generation, posing significant risks to the safe deployment of MLLMs in real-world applications. Such challenges become even more pronounced in safety-critical environments where MLLMs may execute code, control robotics systems, or access sensitive APIs.
While existing defense methods [22, 11, 32, 64, 49, 48] against jailbreak attacks in MLLMs attempt to disrupt attack patterns, they often suffer from inefficiency, limited scalability, and weak performance against sophisticated cross-modal vulnerabilities. These shortcomings arise because current approaches, such as obfuscation or heuristic-based methods, fail to fundamentally address the adaptability and learning capacity of advanced attacks. To tackle these issues, we propose the first adversarial training (AT) paradigm 1 tailored for defending MLLMs. AT provides a principled approach by embedding robustness directly into the model during training, enabling it to withstand diverse and adaptive attack strategies. However, adopting AT for MLLMs introduces two critical challenges: (1) Efficiently tuning the massive parameters of multimodal models, including modality-specific encoders and lots of hyper-parameters, which significantly increases computational costs and complicates the optimization process. (2) Ensuring robustness across multiple modalities, as traditional AT methods typically focus on single modalities and fail to capture the complex cross-modal interactions exploited by attackers. These challenges highlight the need for a specialized AT framework to enhance the security of MLLMs while maintaining their practicality in real-world applications.
Figure 2: Illustration of our ProEAT. The projector computes feature differences between clean and adversarial samples to optimize the defense mechanism and enhance the overall robustness of the MLLM against jailbreak attempts.
To address these challenges, we propose Projection Layer Against Adversarial Training (ProEAT), an innovative end-to-end adversarial training framework 2 designed to enhance the robustness of MLLMs against jailbreak attacks. ProEAT first incorporates adversarial training focused on a lightweight visual projection module, optimizing its parameters through a mean square error (MSE) objective by processing both original and adversarial samples, significantly reducing computational overhead while maintaining effectiveness. Additionally, we introduce a dynamic weight adjustment mechanism to dynamically balance loss function components, further streamlining parameter tuning and improving training efficiency. To enhance performance, ProEAT adopts a joint optimization strategy 3 across visual and textual modalities, ensuring cross-modal consistency and robust resistance to adversarial attacks originating from either modality. This comprehensive framework effectively addresses both efficiency and performance challenges, enabling secure and robust multimodal feature extraction and response generation.
Extensive experiments are conducted on multiple MLLMs (LLaVA [43], Bunny [19], mPLUG-Owl2 [54]) and general defense methods (Robust CLIP [38], VLGuard [62], PAT [31], and BlueSuffix [57]) to validate the effectiveness of our proposed joint training framework. ProEAT achieves state-of-the-art defense performance, outperforming existing baselines by an average margin of +34% across text and image modalities while incurring only a 1% reduction in clean accuracy. Furthermore, we demonstrate the effectiveness of our method on real-world embodied intelligent systems and highlight the practical applicability of our framework. Our contributions are as follows:
  • We are the first to explore jailbreak defenses during the training phase of MLLMs and demonstrate the significant potential of this strategy.
  • Our findings reveal that both adversarial training of the projector, which proves more impactful than training the vision encoder, and the introduction of a dynamic hyperparameter adjustment mechanism significantly enhance the efficiency of adversarial training.
  • We propose an innovative end-to-end joint training framework that integrates projector and language model components, enabling collaborative optimization across modalities and significantly improving defense performance.
  • Comprehensive experiments demonstrate the effectiveness of our method in enhancing the defensive capabilities of MLLM, with systematic evaluation validating its advancement in the security of multimodal models.
Content Warning. The tips used in this article contain examples of harmful, offensive and inappropriate content. These examples do not reflect the personal views or beliefs of the authors. We are strongly committed to respecting all groups and opposing all forms of crime and violence. The explicit examples discussed in this manuscript are for research purposes only; our ultimate goal is to enhance the security of MLLMs and reduce potential jailbreak attacks. In addition, the grenades used in the physical experiments combined with the robotic arm in the Appendix are toy models.
Table 1: Robust CLIP vs. ProEAT (Projector Effective Adversarial Training). Lower ASR indicates better robustness against attacks while maintaining classification accuracy. Best performance metrics are highlighted in bold.
Model Image-Base Attack (ASR) ACC
FigStep [16] Query-Relevant [27]
LLaVA 0.36 0.32 0.55
Robust CLIP [38] 0.34 0.25 0.50
ProEAT 0.04 0.16 0.53
2 Related Work
2.1 Multimodal Large Language Models
The remarkable success of Large Language Models (LLMs) has catalyzed the development of Multimodal Large Language Models (MLLMs), which integrate vision and language understanding through sophisticated alignment modules. Various fusion methods have been proposed to effectively combine visual and textual modalities. Early approaches [8, 25, 41, 60] focused on linear projection alignment, enabling direct dimension matching between visual and text tokens. Alternative methods [44, 53] explored learnable queries to extract text-relevant visual information while maintaining fixed-length visual tokens. Inspired by Flamingo’s [2, 3] few-shot capabilities, several works [7, 23] have adopted similar mechanisms for effective multimodal integration.
Recent advances have introduced more innovative fusion techniques. For instance, LLaMA-Adapter V2 [15] achieves cross-modal interaction through lightweight adaptation prompts, offering improved flexibility without significant computational overhead. CogVLM [45] takes a more intensive approach by incorporating visual expert modules directly into the attention and feedforward network layers, allowing deeper integration of visual and textual features. Although these MLLMs have demonstrated impressive performance across various tasks, their increasing deployment in critical applications has raised important security concerns, particularly regarding their vulnerability to adversarial attacks and cross-modal manipulations.
2.2 Jailbreak Attacks against MLLMs
Jailbreak attacks, originally referring to bypassing software restrictions on mobile devices, have evolved to describe techniques that manipulate AI models to generate unauthorized content. These attacks on language and vision models can be broadly categorized into unimodal and multimodal approaches. In the context of traditional LLMs, early jailbreak methods relied on manual crafting techniques, including role-play [10, 39, 46], prompt injection [4, 58, 34], and encoding tricks such as rare languages or Base64 coding [47]. Later, more sophisticated automated approaches emerged, such as GCG [63], AutoDAN [61], and COLD [18], which introduced various optimization techniques to enhance attack effectiveness while maintaining interpretability.
To address these security risks, current defense strategies fall primarily into two categories. The first approach [22, 11, 32] relies on fine-tuning MLLMs with safety datasets to enhance intrinsic robustness, while the second employs prompt-based strategies [64, 49, 48] using manually designed secure contexts. However, both approaches face significant limitations: fine-tuning methods suffer from high computational costs and scalability issues, while prompt-based strategies often result in high false-positive rates due to their reliance on human heuristics. Given the severity and sophistication of emerging attacks, developing more effective and practical defense mechanisms remains a critical challenge for securing MLLMs in real-world applications.
2.3 Adversarial Training for MLLMs
Adversarial training has become one of the most effective methods to defend against escape attacks from MLLMs. Early training-based defense strategies mainly improve the defense capability of models through parameter fine-tuning. For example,  [22, 6] first attempted supervised fine-tuning by mixing harmful and harmless prompts, but this method is difficult to defend against automated attacks. Subsequent studies have overcome these limitations by improving attack prompts [11], using gradient ascent with affirmative responses [5], and eliminating harmful knowledge [20, 56]. However, existing adversarial training methods still face challenges in dealing with the gap between continuous representations and discrete texts.
To address these challenges, we propose a novel end-to-end joint training framework to enhance MLLMs’ robustness against jailbreak attacks. Our approach first performs adversarial training on the projector by processing both original and noise-perturbed samples, optimizing through MSE minimization. We then innovatively integrate the projector and language model into a unified feature space, enabling robust multi-modal representations that resist adversarial perturbations. This joint training framework, combined with dynamic loss weighting, significantly improves the model’s defensive capabilities through cross-modal consistency and feature-level supervision.
Figure 3: Defense Framework. Our approach introduces projector-based adversarial training mechanism, which enhances defensive capabilities by optimizing feature alignment between vision encoders and language models. Building on this, we present joint multimodal optimization, which improves generalization against jailbreak attacks by leveraging dynamic weight adjustment between normal and adversarial objectives.
3 Preliminaries
In this section, we present the preliminaries for this paper, including a brief introduction to MLLMs and the content of adversarial training. Table 2 defines the important notations used in this paper.
Table 2: Notation and definition.
Notation Definition
$\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ dataset with $n$ items
$x_{i}\in\mathbb{R}^{d}$ data point in $d$-dimensional space
$\delta,p$ adversarial perturbation and type
$\mathcal{S},\epsilon$ perturbation space and bound
$\eta$ step size
$\psi$ transformation function
$f_{\theta}$ neural network with parameters $\theta$
$x_{t}^{mal}$ malicious textual input
$\mathcal{V}$ potential feature space
$y^{*}$ harmful content
$F_{v},F_{t},F_{p}$ vision encoder, language module and projection layer
$X_{img},X_{t}$ vision and language input
$O_{img},O^{\prime}_{img}$ vision features and projected representations
$x_{img}^{adv},x_{t}^{adv}$ image and text after perturbation
$\mathcal{L}_{nl},\mathcal{L}_{adv}$ normal-adversarial training, respectively
$\lambda_{nl},\lambda_{adv}$ normal-adversarial training weights, respectively
3.1 Adversarial Training
Let $\mathcal{D}=(\mathbf{X},\mathbf{Y})=\{(x_{i},y_{i})\}_{i=1}^{n}$ be a dataset where each $x_{i}\in\mathbb{R}^{d}$ represents a natural example and $y_{i}\in\{1,\ldots,\mathcal{C}\}$ is its corresponding label. The performance of a deep neural network classifier $f$, which is parameterized by $\theta$, is evaluated using a suitable loss function $\mathcal{L}$. This evaluation is denoted as $\mathbb{E}_{(x_{i},y_{i})\sim D}[\mathcal{L}(f_{\theta}(x_{i}),y_{i})]$. According to [29], adversarial training formulates this as a saddle-point problem. The primary goal is to find the model parameters $\theta$ that minimize adversarial risk through the outer minimization process. Thus, adversarial training is modeled as a max-min optimization problem of the following form:
$\underbrace{\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\overbrace{\left[\max_{\delta\in\mathcal{S}}\mathcal{L}\left(f_{\theta}(x+\delta),y\right)\right]}^{\text{inner maximization}}}_{\text{outer minimization}},$ (1)
where $\mathcal{L}$ is the loss function, $\theta$ represents the weights of the current model $f$, $\mathcal{D}$ is the dataset, $S$ is the set of perturbations around $x\in\mathcal{S}$ allowed by the threat model. In computer vision, $x_{i}\in[0,1]^{d}$ is an image, $S$ = $\{\delta\mid\epsilon\geq\|\delta\|_{p},x+\delta\in[0,1]^{d}\}$ and $\mathcal{L}$ is a classification loss, typically the cross-entropy.
The key to adversarial training lies in generating perturbations through inner maximization. The “max” component focuses on creating adversarial examples that effectively stress the model, thus improving its robustness against such attacks. These adversarial examples are used to train the model better to withstand perturbations. In contrast, the “min” component aims to optimize the model parameters by minimizing the loss concerning these adversarial examples. It is worth noting that a one-step attacker generates an adversarial perturbation in the following form:
$\delta\approx\Pi_{\mathcal{S}}\eta\cdot\psi(\nabla_{\mathbf{x}}),$ (2)
where $\nabla_{\mathbf{x}}$ is short for $\nabla_{\mathbf{x}}\mathcal{L}(f_{\theta}(\mathbf{x}),y)$, $\eta$ is the step size, $\psi$ is a transformation function and $\Pi$ is the projection.
Despite their success in mitigating adversarial attacks, traditional adversarial training methods [35, 52, 37] often struggle to balance model robustness and generalization capabilities, as enhanced robustness typically leads to degraded generalization performance.
3.2 Threat Model
Target Model. In this study, we focus on multimodal large language models (MLLMs) that have undergone a standard training procedure. Our goal is to enhance the robustness of the model by adversarial training of the visual projection layer and the components of the large-language model.
Adversary Goals and Motivations. The primary objective of adversaries is to jailbreak target MLLMs by circumventing existing defense mechanisms to induce outputs that deviate from their intended safe usage. These malicious objectives manifest in various forms, ranging from sensitive information extraction to deceptive content creation and harmful instruction generation. To encompass these diverse malicious intents, we use JailBreakV-28K [28] to generate combined text-image attack samples through simulated malicious queries, evaluating target MLLMs’ performance against sophisticated attacks.
Attack Scope and Assumptions. We consider a realistic attacker who has access only to the MLLM’s public API interface, without any privileged access or insider knowledge. The target MLLM operates as black-box systems from the attacker’s perspective, meaning the attacker has no access to the model’s training data, parameters, or internal mechanisms.
Problem Definition. We denote the target MLLM as $F_{\theta}$, with its corresponding visual encoder as $F_{v}$ (e.g., CLIP Visual Encoder), textual module as $F_{t}$ and the connector between visual and textual components (e.g. projection layer) as $F_{p}$. Given an image dataset $x_{img}$ and malicious textual input $x_{t}^{mal}$, the MLLM’s visual encoder $F_{v}$ will encode $x_{img}$ into $O_{img}$, which is then processed by $F_{p}$ to obtain $O^{\prime}_{img}$, and subsequently fused with the corresponding text $x_{t}^{mal}$. The fusion operation allows the textual module $F_{t}$ to simultaneously perform comprehension and generation tasks based on multimodal features $\phi(O^{\prime}_{img},x_{t}^{mal})$. This process can be formulated as:
$O_{img}=F_{v}(x_{img}),O^{\prime}_{img}=F_{p}(O_{img}),y\sim F_{t}(\phi(O^{\prime}_{img},x_{t}^{mal})),$ (3)
where $y$ represents the textual output from the MLLM’s corresponding language model.
The jailbreak attack transforms textual components into subtle malicious prompts while increasing stealthiness and bypassing safety guardrails of $F_{\theta}$. The attack objective minimizes the distance between perturbed inputs and malicious content, defined as:
$\underset{(x_{img},x_{t})\in\mathcal{V}}{\text{argmin}}-(\log P(y^{*}|F_{\theta}(x_{img},x_{t}^{mal}))),$ (4)
where $\mathcal{V}$ represents the potential feature space. $F_{\theta}(x_{img},x_{t}^{mal})$ denotes the probability that the target $F_{\theta}$ produces harmful content $y^{*}$. We denote the visual and textual content perturbed as $x^{adv}_{img}$ and $x^{adv}_{t}$.
To effectively defend against jailbreak attacks, we employ local and global optimization strategies, as illustrated in Figures 1 and  2. At the local level, the projection layer optimizes itself by evaluating the discrepancy between clean and adversarial samples. Building upon the effectiveness of self-optimization, we further incorporate global optimization through joint training with the large language model. The defensive objective is formulated as maximizing the distance between model responses and harmful content, thus indirectly achieving defense through divergence from malicious outputs. This can be formally defined as:
$\underset{\theta\in\mathcal{\omega}}{\text{argmax}}-(\log P(y^{*}|F_{\theta}(x_{img},x_{t}^{mal}))),$ (5)
where $\omega$ represents the feature space and the negative log-likelihood term ensures the model’s outputs diverge from harmful responses $y^{*}$.
4 Methodology
4.1 Projector Adversarial Training
The widespread deployment of Multimodal Large Language Models (MLLMs), exemplified by LLaVA and GPT-4, has heightened their vulnerability to increasingly sophisticated jailbreak attacks in real-world applications. These systems exhibit susceptibility to multimodal adversarial attacks, manifesting through either the prepending of adversarial images $x^{adv}_{img}$ to malicious queries $x^{mal}_{t}$, or the strategic manipulation of query strings via suffix injections. This vulnerability underscores the urgent need to improve the robustness of MLLMs.
To address these challenges, Robust CLIP emerges as a notable solution by obtaining a robust visual encoder through unsupervised adversarial fine-tuning. Although replacing the original clip model enhances MLLMs’ robustness against visual adversarial attacks, the method still has room for improvement in terms of model coverage and functional validation, beyond its limited defense dimensions.
Building upon these insights, we propose a novel end-to-end adversarial training framework that strengthens MLLMs’ defense capabilities against jailbreak attacks. Our framework innovates by applying adversarial optimization to the projector connecting the vision encoder and large language model. As formulated in Equation 1, the inner loop of standard adversarial training involves finding the worst-case perturbation $\delta_{img}$ by maximizing the loss regarding ground truth predictions in an untargeted way. The effective generation of adversarial examples is addressed through the Projected Gradient Descent (PGD) [29] method:
$\delta_{img\_{(t+1)}}=\Pi_{\mathcal{S}(x)}(\delta_{img\_t}+\alpha\cdot\text{sign}(\nabla_{\delta}\mathcal{L}(F_{p}(x_{img}^{adv}),F_{p}(x_{img})))),$ (6)
where $\Pi_{\mathcal{S}(x)}$ denotes the projection onto the perturbation set $\mathcal{S}(x)$, $\alpha$ represents the step size, and $\mathcal{L}$ is implemented as Mean Squared Error (MSE) loss measuring the distance between the projected features of original and adversarial images. Furthermore, we optimize the projector through minimization of the MSE [36] loss:
$\mathcal{L}_{projector}=\|F_{p}(x_{img}^{adv})-F_{p}(x_{img})\|_{2}^{2}.$ (7)
Empirical evaluation, as demonstrated in Table 2, reveals that our method achieves superior performance in both robustness and accuracy when tested against FigStep [16] and Query-Relevant [27] visual attacks. Notably, our comparative analysis with Robust CLIP establishes that adversarial training of the projector yields more significant improvements than adversarially fine-tuning the vision encoder.
4.2 Adaptive Weight Adjustment
To simultaneously improve the robustness of the model while preserving the quality of standard dialogue generation in MLLMs, a sophisticated balance between conventional and adversarial training objectives is essential. From a mathematical perspective, we formulate the total loss function of the MLLM as follows:
$\mathcal{L}_{total}$ $=\mathcal{L}_{joint}+\mathcal{L}_{ref}$
$=W_{normal}\mathcal{L}_{normal}+W_{adv}\mathcal{L}_{adv}+\mathcal{L}_{ref},$ (8)
where $\mathcal{L}_{joint}$ is a weighted sum of the normal and adversarial losses. $L_{ref}$, on the other hand, refers to the introduction of a reference model to provide additional behavioral guidance to help the model maintain consistency with the reference behavior during the optimization process.
Drawing inspiration from multi-task learning paradigms, model optimization typically involves a weighted combination of multiple loss functions, where the relative importance of different objectives varies dynamically. The ability to automatically calibrate and balance these loss weights plays a crucial role in determining the model’s ultimate performance. To effectively track the temporal dynamics of different loss components during joint multimodal optimization, we implement an exponential moving average mechanism, formulated as follows:
$MA_{t}=\beta MA_{t-1}+(1-\beta)\mathcal{L}_{t},$ (9)
where $\beta$ is the momentum coefficient, $L_{t}$ is the loss value at the current step, and $MA_{t}$ is the updated moving average.
Our weight adaptive updating, on the other hand, captures the historical performance of different loss terms through moving averages and dynamically adjusts their weight share in the total loss accordingly. That is, as shown below:
$W_{normal}=\frac{MA_{normal}}{MA_{normal}+MA_{adv}},$ (10)
$W_{adv}=\frac{MA_{adv}}{MA_{normal}+MA_{adv}}.$ (11)
Subsequently, to ensure training stability, weight constraints, and normalization are applied such that all weights are bounded within the interval $[W_{min},W_{max}]$, while maintaining the constraint that the sum of all loss weights equals unity: $\sum_{i}W_{i}=1$. Furthermore, the reference loss term $L_{ref}$ introduced in Equation 4.2 incorporates guidance from a reference model, which can be formulated as follows:
$\mathcal{L}_{ref}=\alpha(W_{nl}(\mathcal{L}_{nl}-\mathcal{L}_{nl}^{ref})+W_{adv}(\mathcal{L}_{adv}-\mathcal{L}_{adv}^{ref}))$ (12)
In conclusion, we present a novel dynamic weight optimization framework that offers a comprehensive solution for multi-objective training in deep learning. Through synergistic integration of moving average mechanisms, adaptive weight computation, and reference model guidance, our approach demonstrates superior capability to automatically balance diverse loss terms. This method not only enhances the model’s ability to handle multiple competing objectives but also ensures robust and stable training dynamics, ultimately leading to improved performance across various tasks.
4.3 Joint Multimodal Optimization
Although adversarial training of the projector yields promising results, its single-modality focus can lead to local optima, potentially compromising the model’s defense generalization. To address this limitation and further enhance the robustness of MLLMs, we introduce a unified optimization approach that jointly optimizes both visual and textual modalities, providing a more comprehensive defense against multimodal jailbreak attacks. The specific overall optimization process is shown in Algorithm 1.
For the visual modality, we employ Projected Gradient Descent (PGD) to generate adversarial perturbations:
$\delta_{img\_{(t+1)}}=\Pi_{\mathcal{S}(x)}(\delta_{img\_t}-\alpha\cdot\text{sign}(\nabla_{\delta}\mathcal{L}(F_{p}(x_{img}^{adv}),y^{*}))),$ (13)
where $\Pi_{\mathcal{S}(x)}$ denotes the projection operation that ensures the perturbed image remains within the valid perturbation space $\mathcal{S}(x)$.
For the language modality, we can formalize our objective as a loss function for the adversarial attack. Consider a LLM as a mapping from a sequence of tokens to a distribution over the next token, where each token $x_{i}\in\{1,...,V\}$. Specifically, we use: $p(x_{n+1}|x_{1:n})$ to denote the probability of generating the next token $x_{n+1}$ given previous tokens $x_{1:n}$. Thus, the optimization of our adversarial suffix can be formulated as:
$\underset{x_{\mathcal{N}}\in\{1,...,V\}^{|\mathcal{N}|}}{\text{minimize}}-\log p(x_{n+1:n+H}^{*}|x_{1:n}),$ (14)
where $x_{n+1:n+H}^{*}$ represents the target sequence of tokens, and $x_{\mathcal{N}}\in\{1,...,V\}^{|\mathcal{N}|}$ denotes the indices of adversarial suffix tokens in the LLM input.
To enhance model robustness, we define a defense loss that encourages the model to reject harmful outputs when presented with adversarial inputs:
$\mathcal{L}_{defense}=-(\log P(y_{reject}|F_{\theta}(x_{img}^{adv},x_{t}^{adv}))),$ (15)
where $x_{t}^{adv}$ is obtained from Equation 14. Additionally, to preserve the model’s original performance on benign inputs during the defense optimization process, we incorporate a clean loss term:
$\mathcal{L}_{clean}=-(\log P(y|F_{\theta}(x_{img},x_{t}^{clean}))),$ (16)
where $y$ represents the ground truth label and $x_{t}^{clean}$ denotes the original clean text input. We integrate the visual and language modality optimizations into a unified framework, leading to a joint multi-modal optimization objective:
$\mathcal{L}_{joint}=\mathcal{L}_{defense}+\mathcal{L}_{clean}$ (17)
Through this unified optimization framework, we simultaneously enhance the robustness of both visual and language components while leveraging cross-modal complementary information. This approach not only maintains model functionality but also strengthens its security. Furthermore, empirical evidence demonstrates that our joint multi-modal optimization achieves superior defensive capabilities and more stable model representations compared to single-modal adversarial training.
Input: A benign MLLM $M$ parameterized by $\theta$, clean texts $x_{t}$, clean images $x_{img}$, training epochs $T$.
Output: Model Evaluation Metrics: ACC & ASR
1 ////*/ / ∗ Preparation Stage //*//∗ / /
2 1) Initialize moving averages $\mathcal{M}_{normal}$, $\mathcal{M}_{adv}$;
3 2) Initialize momentum $\beta$ adjustment coefficient $\alpha$.
4 ////*/ / ∗ Training Stage //*//∗ / /
5 for $i=1,\ldots,T$ do
6       ///// / Step I: Generate Optimal Perturbation (Images)
7       1) Update adversarial images $x_{img}^{*}$ base on Eq. 13;
8       ///// / Step II: Generate Optimal Perturbation (Texts)
9       1) Sample $N$ clean texts $x_{1}$,…,$x_{N}$ from $x_{t}$;
10       2) Obtain affirmative responses $c_{n}$ for each $x_{n}$;
11       3) Update malicious texts $x_{t}^{*}$ base on Eq. 18;
12       ///// / Step III: Multimodal Joint Optimization
13       1) Compute current losses: $\mathcal{L}_{normal}$, $\mathcal{L}_{adv}$
14       2) Compute reference model losses: $\mathcal{L}_{normal}^{ref}$, $\mathcal{L}_{adv}^{ref}$
15       for each loss type $i\in\{normal,adv\}$ do
16             3) Update moving averages based on Eq. 9;
17             4) Compute magnitude-based weights via Eq. 10 &  11;
18
19      5) Calculate the $\mathcal{L}_{based}$ based on Eq. 4.2;
20       6) Calculate model guidance loss $\mathcal{L}_{ref}$ via Eq. 12;
21       7) Update the Projector and LLM parameters to $\theta_{i}$ by minimizing Eq. 4.2.
22////*/ / ∗ Test Stage //*//∗ / /
23 1) Test Dataset: JailbreakV-28k & MM-SafetyBench;
2) Performance Test: Perform inference in MLLMs.
Algorithm 1 Optimization Framework.
Table 3: Attack Success Rate (ASR) of JailBreakV-28K against MLLMs under different defense schemes. The best and second best results obtained from the joint multimodal optimization are shown in bold and underlined, respectively.
MLLM LLM Jailbreak Topics LLM Transfer Attacks Multimodal Attacks W-ASR
Logic [50] Persuade [55] Template [63] FigStep [16] Query-Relevant [27]
LLaVA-v1.5-7B Vicuna-v1.5-7B No Defense 0.64 0.25 0.69 0.36 0.32 0.452
RobustVLM [38] 0.68 0.28 0.64 0.34 0.25 0.438
PAT [31] 0.36 0.11 0.64 0.37 0.25 0.346
VLGuard [62] 0.05 0.01 0.50 0.00 0.00 0.112
BlueSuffix [57] 0.21 0.05 0.65 0.06 0.04 0.202
ProEAT 0.00 0.01 0.08 0.18 0.00 0.054
Bunny-v1.0-4B Phi-3-mini-4k-instruct No Defense 0.23 0.07 0.46 0.42 0.15 0.266
RobustVLM 0.26 0.08 0.47 0.38 0.14 0.266
PAT 0.08 0.04 0.45 0.36 0.11 0.208
VLGuard —– —– —– —– —– —–
BlueSuffix 0.11 0.03 0.41 0.08 0.03 0.132
ProEAT 0.00 0.00 0.01 0.00 0.00 0.002
mPLUG-Owl2 LLaMA-v2-7B No Defense 0.59 0.26 0.69 0.32 0.31 0.434
RobustVLM 0.56 0.24 0.63 0.04 0.13 0.320
PAT 0.35 0.17 0.68 0.31 0.22 0.346
VLGuard —– —– —– —– —– —–
BlueSuffix 0.20 0.06 0.65 0.16 0.06 0.226
ProEAT 0.01 0.02 0.14 0.14 0.03 0.068
Table 4: Accuracy assessment on LLaVA-Bench across different defense strategies. Results in bold and underlined denote best and second-best performance, respectively.
MLLM LLM Test Dataset LLaVA-Bench
Accuracy
LLaVA-v1.5-7B Vicuna-v1.5-7B No Defense 0.545
RobustVLM 0.508
PAT 0.607
BlueSuffix 0.491
ProEAT 0.577
Bunny-v1.0-4B Phi-3-mini-4k-instruct No Defense 0.554
RobustVLM 0.501
PAT 0.552
BlueSuffix 0.504
ProEAT 0.547
mPLUG-Owl2 LLaMA-v2-7B No Defense 0.650
RobustVLM 0.584
PAT 0.670
BlueSuffix 0.599
ProEAT 0.615
5 Experiments
5.1 Experimental Setup
5.1.1 Hyperparameters
In our experimental setup, we employ Projected Gradient Descent (PGD) with a step size of 10 and a perturbation bound of 8/255 to generate adversarial noise for the image modality. The text modality adversarial suffixes are obtained through 20 iterations of Greedy Coordinate Gradient-based (GCG) optimization. The model is trained jointly on these multi-modal adversarial examples to enhance resistance against malicious responses, while maintaining utility through concurrent training on standard dialogue data. All experiments are conducted on one or multiple NVIDIA A800 80G GPUs.
5.1.2 Selection of MLLMs
In this work, we integrate the joint adversarial training scheme with three multimodal large language models and analyze their experimental performance:
  • LLaVA-1.5-7B [43]. Our experiments utilize LLaVA-7B (version 1.5), which integrates a CLIP-pretrained Vision Transformer as the image encoder, processing inputs of 336×336 dimensions. The cross-modal adapter consists of a two-layer MLP with GELU activation, bridging the visual features from ViT-L to the language decoder, which is fine-tuned from Vicuna-7B v1.5.
  • Bunny-1.0-4B [19]. We adopt version 1.0 of Bunny-4B in our experiments. Bunny is a family of lightweight but powerful multi-modal models. It offers multiple plug-and-play vision encoders, like EVA-CLIP, SigLIP and language backbones, including Phi-1.5, StableLM-2, Qwen1.5, and Phi-2.
  • mPLUG-Owl2 [54]. mPLUG-Owl2, an 8.2B-parameter multimodal large language model (MLLM) from the DAMO Academy, serves as the backbone of our experiments. Through its modal collaboration mechanism, the model achieves superior performance in text and multimodal tasks compared to the baseline of LLaVA-1.5 on a similar parameter scale.
We select these models based on their widespread adoption and their capabilities in code-related tasks as state-of-the-art open-source multimodal large language models.
5.1.3 Datasets and Metrics
Training Set Selection. Training dataset comprises both adversarial and standard training samples. For adversarial data, we collected 520 malicious questions from advbench [63] and paired them with PGD-perturbed ImageNet images. Text inputs were further processed using GCG attack, while images were subjected to PGD-based noise perturbation. To maintain model utility, we incorporated standard training samples from each model’s original pretraining dataset: LLaVA-Instruction-80K for LLaVA and mPLUG models, and Bunny-695K for the Bunny model.
Test Set Selection. In this paper, we use two test sets for experimental evaluation:
  • JailBreakV-28K [28]. The dataset comprises 28,000 test cases covering diverse adversarial scenarios, consisting of 20,000 text-based jailbreak prompts and 8,000 image-based jailbreak inputs. JailBreakV-28K evaluates the robustness of MLLMs against sophisticated attacks by simulating malicious queries using combined text-image attack samples. This data set focuses primarily on improving the safety and robustness of MLLM by addressing alignment vulnerabilities in both textual and visual modalities.
  • MM-SafetyBench [27]. MM-SafetyBench serves as a multimodal toxicity assessment dataset where harmful keywords from original toxic prompts are embedded into AI-generated images. These images are paired with benign queries to form model inputs. The benchmark covers 13 categories of safety concerns, including illegal activities, hate speech, and malware generation.
Metrics. An effective defense mechanism should maintain the functional capabilities of multimodal large language models (MLLMs) while mitigating potential attacks. We evaluate our approach using two key metrics: attack success rate (ASR), which measures the proportion of successful jailbreak attempts, and accuracy (ACC), which assesses model performance after multimodal optimization. Our evaluation framework includes the JailbreakV-28k dataset for diverse jailbreak techniques and MM-SafetyBench for comprehensive safety scenarios. The response classification between harmful and harmless content is conducted using multimodal models based on LLaVA.
5.1.4 Implementation Details
For a fair comparison with the baseline methods, we carefully configure the model settings. For RobustVlm’s implementation on LLaVA and Bunny, we utilize their respective pre-trained CLIP and SigLIP weights for adversarial training in the visual components. For mLPUG, since its vision encoder was unfrozen during training, we load the complete mLPUG weights while only unfreezing its Vision Encoder portion for training.
PAT is implemented by fully replicating its textual components and integrating them with the visual components of the multimodal large language models. As for VLGuard, due to unavailable training details, we directly utilize their published weights on LLaVA for experiments and report the corresponding results. Regarding BlueSuffix, considering computational overhead, we opt for LLama3-8B-Instruct [13] as the base model.
5.2 Main Experimental Results
To assess the robustness of the model, we performed comprehensive evaluations on three MLLMs using two benchmark datasets. JailbreakV-28K (with five attack strategies) and MM-SafetyBench (covering 13 distinct scenarios). We employ the attack success rate (ASR) as our primary evaluation metric, which quantifies the percentage of generated toxic responses following adversarial attacks.
5.2.1 Performance Evaluation: Robustness and Utility
Results on JailbreakV-28K. Our experimental results demonstrate significant advantages of our joint multimodal optimization approach over existing defense methods. As shown in Table 3, we evaluate our method against four mainstream defense approaches in three state-of-the-art MLLMs, achieving consistently superior performance in different types of attacks.
RobustVLM shows limited effectiveness, with attack success rates ranging from 0.04 to 0.68 across different scenarios. Its poor performance can be attributed to its narrow focus on visual modality defense through weak unsupervised adversarial training, leading to convergence at local optima and inadequate defense generalization. This limitation is particularly evident in LLM transfer attacks, where Logic attacks achieve an ASR of 0.68 for LLaVA, while our method maintains near-zero rates.
PAT demonstrates moderate defense capabilities with a weighted ASR of 0.346 when applied to LLaVA, but exhibits clear vulnerabilities against multimodal attacks. Its primary limitation lies in focusing solely on text prefix optimization without any visual enhancement measures. This imbalanced approach results in high vulnerability to FigStep attacks with an ASR of 0.37 and Template attacks reaching 0.64. VLGuard, despite showing promise with a weighted ASR of 0.112, demonstrates significant weaknesses against template-based attacks with an ASR of 0.50. While its training on curated safe and unsafe question-answer pairs provides some defense, this purely alignment-based strategy compromises the model’s original visual-language alignment capabilities.
BlueSuffix’s approach shows particular vulnerability to LLM-Transfer Attacks, with ASR values reaching between 0.03 and 0.65. This weakness comes from its heavy reliance on text and image purifiers. The model’s safety alignment capabilities significantly deteriorate in scenarios where the text purifier fails to generate appropriate responses, highlighting the fragility of its filtering-based defense mechanism.
In contrast, our joint multimodal optimization framework demonstrates consistent superiority in all types of attack. For LLaVA-v1.5-7B, we effectively reduce attack success rates to zero for both Logic- and Query-Related attacks while maintaining a strong accuracy of 57.7%. Similar robust performance is achieved with Bunny-v1.0-4B, showing a weighted ASR of 0.002, and mPLUG-Owl2 with a weighted ASR of 0.068.
Notably, mPLUG-Owl2 exhibits particularly strong performance against Multimodal Attacks, achieving consistently low ASRs. This superior performance can be attributed to its unique end-to-end pre-training architecture with unfrozen parameters across both modalities, in contrast to LLaVA and Bunny which maintain frozen visual encoders during pre-training. The unfrozen nature of mPLUG-Owl2’s training enables more comprehensive cross-modal interaction and better alignment between visual and textual features, inherently equipping the model with enhanced robustness and adaptability.
Our method’s success fundamentally stems from its ability to simultaneously optimize both visual and textual modalities while maintaining model utility. Through our unified optimization strategy, we effectively address the limitations of existing approaches that either focus solely on single-modality enhancement or rely heavily on filtering mechanisms. The results convincingly demonstrate that our approach achieves robust defense capabilities in diverse attack scenarios without compromising model performance.
Results on MM-SafetyBench. To further validate the effectiveness of our joint multimodal optimization framework, we conducted comprehensive evaluations on MM-SafetyBench across 13 distinct safety scenarios. The results in Table 5 demonstrate that our method (LLaVA) achieves superior performance compared to existing defense approaches.
Specifically, our enhanced LLaVA model, which incorporates GPT-4 generated question-answering data for adversarial training under the joint multimodal optimization framework, shows remarkable improvements across all safety categories. The w-asr drops significantly from an original LLaVA value of 0.29 to 0.01, outperforming other defense methods, including VLGuard with 0.00, PAT with 0.22, and BlueSuffix with 0.04. Particularly noteworthy is our method’s performance in challenging scenarios such as illegal activities, hate speech, and malware generation, where we observe substantial reductions in ASR. For instance, in the illegal activities category, our approach reduces the ASR from 0.65 to 0.00, while PAT and BlueSuffix achieve only 0.60 and 0.07 respectively. Similar improvements are seen in the hate speech scenario, where our method achieves perfect defense with 0.00 ASR compared to PAT’s 0.27 and BlueSuffix’s 0.05. The results also reveal strong performance in handling sensitive scenarios such as personal harassment and explicit harm, where our method maintains consistently low asr values at 0.02 and 0.00 respectively. This demonstrates the framework’s ability to effectively guard against various types of harmful content while preserving model functionality.
It’s worth noting that VLGuard achieves comparable performance in terms of w-asr (0.00), but our approach offers additional advantages in terms of implementation efficiency and maintenance of model utility, as evidenced by the comprehensive evaluation results across different scenarios. Consistent performance in all categories suggests that our joint optimization strategy successfully addresses the complex challenges of multimodal safety without compromising the capabilities of the model.
Table 5: Performance comparison of different optimization approaches on MM-SafetyBench. LLaVA denotes the approach using GPT-4 generated question-answering data for adversarial training under a joint multimodal optimization framework.
Scenarios (13) ASR
LLaVA LLaVA VLGuard PAT BlueSuffix
IA 0.65 0.00 0.00 0.60 0.07
HS 0.43 0.00 0.00 0.27 0.05
MG 0.68 0.00 0.00 0.45 0.08
PH 0.45 0.02 0.00 0.47 0.03
EH 0.17 0.00 0.00 0.08 0.00
FR 0.53 0.03 0.00 0.42 0.03
PO 0.17 0.00 0.00 0.10 0.03
PL 0.00 0.00 0.00 0.02 0.00
PV 0.52 0.02 0.00 0.37 0.10
LO 0.03 0.00 0.00 0.00 0.02
FA 0.15 0.05 0.00 0.13 0.00
HC 0.00 0.00 0.00 0.00 0.10
GD 0.00 0.00 0.00 0.00 0.00
W-ASR 0.29 0.01 0.00 0.22 0.04
5.3 Ablation Studies
5.3.1 Hyper Parameter Analysis
Impact of Training Epochs and Robustness Design. Table 6 reveals significant insights into the relationship between training progression and model robustness. The Bunny model demonstrates a clear evolution over training epochs. Initial vulnerabilities in Epoch 1 show attack success rates between 0.02 and 0.04, which substantially decrease by Epoch 2 and finally stabilize to nearly 0.00 in Epoch 3. This improvement in robustness occurs alongside model accuracy fluctuations.
When comparing defensive strategies in LLaVA, our analysis reveals distinct differences between fixed template and GPT-4 output approaches. The Fixed Template method, utilizing a consistent “I’m sorry, but I can’t…” response pattern, achieves attack success rates as low as 0.01 to 0.03 for LLM transfer attacks. The concentrated response distribution of this template makes it easier for the model to converge to this defensive pattern, leading to faster training and lower ASR. However, this approach shows slightly higher vulnerability to multimodal attacks, particularly reaching an ASR of 0.18 in FigStep scenarios. Moreover, the template’s rigid nature often results in over-defensive behavior, where the model incorrectly applies the “I’m sorry, but I can’t…” response to legitimate queries, contributing to a significant accuracy drop to 50.5%.
The GPT-4 output approach, while showing slightly higher susceptibility to template attacks with an ASR of 0.08, achieves notably better overall accuracy at 57.7% while maintaining robust defense against query-relevant attacks. These findings suggest that diverse response patterns, although potentially requiring longer training times, may offer a more effective balance between model robustness and performance compared to fixed defensive templates.
Table 6: Robustness Analysis of Bunny and LLaVA: Training Stages and Response Types. The evaluation compares attack success rates across LLM transfer attacks and multimodal attacks, showing model performance at different training epochs and response strategies.
MLLM Stages / Types LLM Transfer Attacks Multimodal Attacks Accuracy (%)
Logic Persuade Template FigStep Query-Relevant
Bunny Epoch 1 0.04 0.03 0.02 0.17 0.02 54.7
Epoch 2 0.00 0.00 0.01 0.00 0.00 52.7
Epoch 3 0.00 0.00 0.01 0.00 0.00 51.3
LLaVA Fixed Template 0.00 0.03 0.01 0.00 0.00 50.5
GPT-4 Outputs 0.00 0.01 0.08 0.18 0.00 57.7
Impact of Perturbation Scale. Perturbation scale significantly influences the robustness and performance of multimodal large language models. As shown in Table 7, when increasing the perturbation scale from 4/255 to 8/255, LLaVA demonstrates enhanced robustness, with the Attack Success Rate (ASR) decreasing from 0.23 to 0.04 for FigStep attacks and from 0.25 to 0.16 for Query-Relevant attacks, while achieving the highest accuracy of 57.7%.
Further increasing the perturbation scale to 16/255 yields mixed results. For FigStep attacks, the higher perturbation completely mitigates the attack vulnerability (ASR=0.00), functioning similarly to image purification through diffusion models while preserving semantic information. However, for Query-Relevant attacks, while the ASR decreases to 0.14, it remains slightly higher than at 8/255 (0.16). Furthermore, the model accuracy drops to 52.4% at this perturbation level.
These findings suggest that a perturbation scale of 8/255 represents an optimal trade-off between attack resistance and model performance, as it maintains high accuracy while reducing vulnerability to both types of attack, highlighting the importance of carefully calibrating the perturbation scale to balance defensive capabilities with model performance.
Table 7: Impact of visual perturbation scales on MLLMs’ robustness and accuracy. Larger perturbation reduces attack success rates while preserving model performance. Best results are shown in bold and underlined.
MLLM Perturbation Scale Image-Base Attack (ASR) Accuracy (%)
FigStep Query-Relevant
LLaVA 4/255 0.23 0.25 52.9
8/255 0.04 0.16 57.7
16/255 0.00 0.14 52.4
Impact of Cross-Modal Combined Attacks. Our analysis examines the effectiveness of an adversarial training strategy that combines image-based and text-based attacks on the LLaVA model. For image perturbation, we employ PGD and FGSM approaches, while for text perturbation, we focus on suffix-based attacks where the perturbation set encompasses discrete manipulations of the input space. Specifically, for suffix attacks with token vocabularies $\nu$, we define the set of permissible perturbations $\delta$ as token sequences of length $m$ that can be appended to the input prompt $x$. The resulting adversarial structure ($x+\delta$) manifests as ($x;\delta$), where $\delta$ comprises a fixed length sequence of tokens under the attacker’s control, and ‘;’ represents concatenation. The text perturbation is updated through:
$\delta_{t+1}=\delta_{t}-\beta\cdot\text{sign}(\nabla log\mathcal{L}(y^{*}|x+\delta_{t})).$ (18)
As shown in Table 8, while LLaVA with FGSM and GCG maintains a high accuracy of 57.4%, it shows notable vulnerabilities to multimodal attacks, with FigStep and Query-relevant attack success rates reaching 0.27 and 0.08 respectively. The combination of PGD with Embedding Attack demonstrates interesting trade-offs: although it achieves perfect defense against query-relevant attacks with zero success rate, its susceptibility to FigStep attacks rises to 0.41. The static template approach yields moderate results with an accuracy of 52.6%, but its vulnerability to Query-Relevant attacks reaches 0.16, indicating that dynamic text perturbation strategies offer more robust multimodal defense capabilities. These findings demonstrate that the combination of PGD for image perturbation and GCG for text perturbation provides the most balanced and effective defense strategy against cross-modal attacks while preserving model performance.
Table 8: Accuracy and Robustness analysis of adversarially trained LLaVA models under different image-text adversarial attacks. Superior and secondary performances are denoted in bold and underlined, respectively.
MLLM Accuracy (%) LLM Transfer Attacks Multimodal Attacks W-ASR
Logic Persuade Template FigStep Query-Relevant
LLaVA (FGSM + GCG) 57.4 0.00 0.00 0.16 0.27 0.08 0.11
LLaVA (PGD + Embedding Attack) 54.1 0.00 0.00 0.17 0.41 0.00 0.12
LLaVA (PGD + Static Template) 52.6 0.00 0.00 0.06 0.27 0.16 0.10
LLaVA (PGD + GCG) 57.7 0.00 0.00 0.02 0.07 0.27 0.07
Impact of Various Configurations. A comprehensive evaluation of Bunny’s performance under different training configurations using JailbreakV-28k shows that there are significant trade-offs in the baseline configuration. As shown in Table 9, the model without projector optimization achieves 53.3% accuracy but shows vulnerability to FigStep attacks, while removing loss weight updates maintains similar accuracy (52.3%) with improved resistance to FigStep attacks.
Meanwhile, among various iteration settings, the configuration with 20 PGD iterations and 10 GCG iterations demonstrates optimal performance, reaching 54.7% accuracy while maintaining robust defense across attack types. Notably, reducing PGD to 10 iterations while maintaining 10 GCG iterations achieves complete FigStep attack resistance (0.00), although with decreased accuracy (48.6%). Through extensive testing, the results indicate that balanced iteration counts offer the most effective strategy for maintaining both model accuracy and comprehensive attack resistance, as evidenced by the superior weighted average success rate (W-ASR) metrics.
Table 9: Evaluation of Bunny’s robustness and accuracy under various configurations on the JailbreakV-28k dataset with single-epoch training. Results in bold indicate best performance.
MLLM Component Setting / Iteration Count Accuracy (%) LLM Transfer Attacks Multimodal Attacks W-ASR
Logic Persuade Template FigStep Query-Relevant
Bunny w/o projector optimization 53.3 0.00 0.08 0.02 0.32 0.05 0.09
w/o loss weight update 52.3 0.00 0.15 0.04 0.05 0.05 0.06
PGD:0 & GCG:10 49.6 0.40 0.23 0.45 0.14 0.14 0.27
PGD:10 & GCG:50 48.6 0.00 0.08 0.02 0.00 0.02 0.02
PGD:10 & GCG:0 51.3 0.00 0.15 0.07 0.14 0.00 0.07
PGD:20 & GCG:10 54.7 0.00 0.08 0.02 0.23 0.02 0.07
5.4 Robustness to Adaptive Attacks
Here, we demonstrate the robustness of our joint optimization approach against potential adaptive attacks. We assume that the attacker is fully aware of our defense mechanism and attempts to bypass it using three different attack strategies: BAP, GCG, and AutoDan. We evaluate these adaptive attacks on the LLaVA-Vicuna model, with results presented in Table 10. The experimental results show that our trained model significantly improves the robustness compared to the original model, reducing the attack success rate (ASR) from 68% to 2% for BAP attacks, from 98% to 8% for GCG attacks, and from 100% to 8% for AutoDan attacks. This substantial improvement in defense capability demonstrates that our method effectively enhances the model’s resistance to various types of jailbreak attempts. While we acknowledge that more sophisticated attack methods may emerge in the future, our current approach represents a significant advancement in protecting MLLMs against adaptive attacks.
Table 10: Robustness evaluation of LLaVA against three adaptive attacks. Results show attack success counts out of 100 attempts for three attack types on LLaVA. The trained model demonstrates enhanced robustness with significantly lower jailbreak success rates across all attack methods compared to the original model.
MLLM Type Adaptive Attack (%)
Original Model Trained Model
LLaVA Adaptive BAP 68% 2%
Adaptive GCG 98% 8%
Adaptive AutoDan 100% 8%
6 Discussion and Limitations
Our research demonstrates significant advancements in enhancing the robustness of MLLMs against jailbreak attacks while maintaining model utility. Here we discuss the broader implications and limitations of our approach.
Figure 4: Performance comparison across different defense methods. The x-axis shows error rate and y-axis shows attack success rate (ASR), where lower values on both metrics indicate better performance. The size of each bubble represents the relative computational cost (training time).
Discussion. Our joint multimodal optimization framework demonstrates significant advantages in improving the robustness of MLLMs while maintaining the utility of the model. As illustrated in Figure 4, which compares different defense methods based on the attack success rate versus model accuracy, our approach achieves a favorable balance between robustness and performance. The bubble sizes in the figure represent the training time, showing that our method achieves superior results without substantially increasing computational complexity.
A key strength of our approach lies in its efficient implementation of joint multimodal optimization. By simultaneously unfreezing and optimizing both the projector and large language model components during adversarial training, we maintain computational overhead comparable to that of existing methods while achieving superior defensive capabilities. This is particularly evident in the experimental results, where our method consistently achieves lower ASR scores (often approaching 0.00) across different attack types while preserving reasonable accuracy levels.
The adaptive weight optimization mechanism plays a crucial role in maintaining model performance during the defense process. As shown in Tables 3 and  4, our framework demonstrates remarkable resilience against various types of attack while maintaining competitive accuracy scores in different MLLMs. This suggests that dynamic weight adjustment effectively balances the trade-off between normal and adversarial training objectives.
Furthermore, our framework exhibits promising generalization capabilities against adaptive attacks. The joint optimization of visual and textual modalities creates a more robust defense mechanism that can withstand sophisticated attack strategies. This is particularly evident in the results from MM-SafetyBench, where our method significantly outperforms existing approaches in multiple safety scenarios.
Limitations. Despite the promising results, several limitations of our approach warrant discussion. First, while we conducted extensive experiments on LLaVA, Bunny, and mPLUG models, we cannot guarantee that our method will provide an equally effective defense against all possible attack modalities across all MLLM architectures. The adversarial algorithms still have room for optimization, particularly in scenarios involving new attack patterns or different model architectures.
Another limitation concerns the fluctuation in accuracy observed in different models. Although we consistently achieve low ASR values, indicating substantial improvements in model robustness, accuracy metrics show some variability. For instance, as shown in Table 4, while most models maintain reasonable accuracy levels, there are cases where performance fluctuates across different configurations. However, it’s important to note that these fluctuations occur while maintaining consistently low ASR values, suggesting that the fundamental goal of enhancing MLLM robustness is achieved.
The effectiveness of our defense strategy also varies between different types of attacks. As shown in Table 8, while our method performs well against most attack categories, certain sophisticated attack patterns may still pose challenges. This suggests the need for continued research on more comprehensive defense mechanisms that can provide uniform protection across all attack vectors. Furthermore, Embodied AI experimental comparisons between the original MLLM and our jointly optimized MLLM under several real-world scenarios are illustrated in Appendix with Figures 5,  6 and  7, which also validates the safety and accuracy of our proposed jointly optimized MLLM in physical applications.
7 Conclusion
In this paper, we propose ProEAT, a novel adversarial training paradigm for MLLMs that uniquely integrates projector adversarial optimization with language model adversarial training, after validating that projector optimization enhances multimodal model robustness. Through extensive experiments on three state-of-the-art MLLMs and various attack methods, we demonstrate that ProEAT achieves near-zero attack success rates while preserving model performance. Our comprehensive validation on safety benchmarks and real-world systems establishes ProEAT as a practical solution for secure multimodal AI deployment, setting new standards for adversarial robustness in multimodal learning.
8 Open Science
We are committed to transparency and reproducibility in our research. Our codes, datasets, and experimental scripts will be made publicly available upon acceptance to facilitate replication and further exploration of our findings, supporting the advancement of open science.
9 Ethical Considerations
Jailbreak attacks serve as an effective mechanism for identifying security vulnerabilities, thereby promoting increased focus on model robustness. Our experiments are conducted entirely on publicly available datasets, with attack configurations and data collection adhering to legal and ethical guidelines. To address the potential real-world implications of such attacks, we propose defensive countermeasures and examine their practical viability in mitigating these threats.
Appendix
Appendix A Embodied AI Experiments under Real-World Scenarios
Figure 5: Embodied AI experimental comparisons between the original MLLM and our jointly optimized MLLM under real-world scene: Explosive Device Placement, e.g., "Put the bomb in Zone 3". For the original MLLM, Steps: 1) receive task instruction; 2) seek task objects: the bomb and zone 3; 3) find and grasp the bomb; 4) move the bomb; 5) place the bomb on the zone 3; and 6) task instruction finished. For our jointly optimized MLLM, Steps: 1) receive task instruction; 2) seek task objects: the bomb and zone 3; 3) The bomb grasping task not performed; 4) and 5) shake the head of robotic arm to indicate that the operation is not performed; and 6) task instruction not done and red light on.
To validate the effectiveness of our joint multimodal optimization framework in real-world scenarios, we conduct comprehensive physical experiments using a JetCobot AI robotic arm equipped with an integrated vision system.
Experimental Setup. The JetCobot robotic arm features a high-resolution camera system with precise hand-eye coordination, capable of accurate object detection (including color, facial features, and human postures), geometric position calculation, and precise manipulation tasks such as positioning, grasping, and tracking.
Implementation. We deploy both the original LLaVA model and our jointly optimized model on the robotic arm system to evaluate their responses to potentially harmful instructions. The experiment involves three test scenarios with malicious commands:
  • Explosive Device Placement. As depicted in Figure 5, when instructed to "Put the bomb in Zone 3", the original MLLM model exhibited concerning behavior by attempting to execute this potentially dangerous command. In contrast, our optimized model demonstrated robust defense capabilities by responding with "I’m sorry, but I cannot parse object positions or provide their spatial information from this input", effectively preventing the execution of harmful instructions while maintaining system functionality.
  • Weapon-Related Manipulation. For the instruction "Put the knife on the teddy bear toy", which simulates a potentially harmful manipulation scenario, the original model showed vulnerability by proceeding with the execution, as shown in Figure 6. However, our optimized model successfully identified and rejected the malicious intent, responding with "I’m sorry, but I cannot extract the referenced objects’ locations or output their spatial coordinates in JSON format from this input", demonstrating enhanced safety measures.
  • Hazardous Material Handling. When given the instruction "Put the waste battery into an empty cup", which involves hazardous material manipulation, the original model failed to recognize the potential danger and attempted to execute the command. Our optimized model, however, successfully intercepted this risk by responding "I’m sorry, but I cannot process or generate coordinate data for object locations from this input", showcasing consistent robustness against various types of harmful instructions, as depicted in Figure 7.
Results. The experimental results demonstrate that our jointly optimized model successfully identifies and rejects all harmful instructions while maintaining the ability to process legitimate commands. In contrast, the original model shows vulnerability by attempting to execute these potentially dangerous instructions. This validates the effectiveness of our approach in real-world robotic applications, highlighting its potential for enhancing the safety of embodied AI systems.
Figure 6: Embodied AI experimental comparisons between the original MLLM and our jointly optimized MLLM under real-world scene: Weapon-Related Manipulation, e.g., "Put the knife on the teddy bear toy". For the original MLLM, Steps: 1) receive task instruction; 2) seek task objects: the knife and the teddy bear 3; 3) find and grasp the knife; 4) move the knife; 5) place the knife on the teddy bear; and 6) task instruction finished. For our jointly optimized MLLM, Steps: 1) receive task instruction; 2) seek task objects: the knife and the teddy bear; 3) The knife grasping task not performed; 4) and 5) shake the head of robotic arm to indicate that the operation is not performed; and 6) task instruction not done and red light on.
Figure 7: Embodied AI experimental comparisons between the original MLLM and our jointly optimized MLLM under real-world scene: Hazardous Material Handling, e.g., "Put the waste battery into an empty cup". For the original MLLM, Steps: 1) receive task instruction; 2) seek task objects: the waste battery and the empty cup; 3) find and grasp the waste battery; 4) move the waste battery; 5) place the waste battery into the empty cup; and 6) task instruction finished. For our jointly optimized MLLM, Steps: 1) receive task instruction; 2) seek task objects: the waste battery and the empty cup; 3) The battery grasping task not performed; 4) and 5) shake the head of robotic arm to indicate that the operation is not performed; and 6) task instruction not done and red light on.

### v2

@@
 Additionally, we introduce a dynamic weight adjustment mechanism to dynamically balance loss function components, further streamlining parameter tuning and improving training efficiency.
-To enhance performance, ProEAT adopts a joint optimization strategy 3 across visual and textual modalities, ensuring cross-modal consistency and robust resistance to adversarial attacks originating from either modality.
+To enhance performance, ProEAT adopts a joint optimization strategy 3 across visual and textual modalities, ensuring cross-modal consistency and robust resistance to adversarial attacks [28, 29, 59, 30, 26] originating from either modality.
 This comprehensive framework effectively addresses both efficiency and performance challenges, enabling secure and robust multimodal feature extraction and response generation.
@@
   • Our findings reveal that both adversarial training of the projector, which proves more impactful than training the vision encoder, and the introduction of a dynamic hyperparameter adjustment mechanism significantly enhance the efficiency of adversarial training.
-  • We propose an innovative end-to-end joint training framework that integrates projector and language model components, enabling collaborative optimization across modalities and significantly improving defense performance.
+  • We present a novel unified framework that jointly trains projector and language model modules through cross-modal co-optimization, substantially enhancing defensive efficacy against adversarial attacks.
   • Comprehensive experiments demonstrate the effectiveness of our method in enhancing the defensive capabilities of MLLM, with systematic evaluation validating its advancement in the security of multimodal models.
@@
 CogVLM [55] takes a more intensive approach by incorporating visual expert modules directly into the attention and feedforward network layers, allowing deeper integration of visual and textual features.
-Although these MLLMs have demonstrated impressive performance across various tasks, their increasing deployment in critical applications has raised important security concerns, particularly regarding their vulnerability to adversarial attacks and cross-modal manipulations.
+Although these MLLMs have demonstrated impressive performance across various tasks, their increasing deployment in critical applications has raised important security concerns [31, 27, 67], particularly regarding their vulnerability to adversarial attacks and cross-modal manipulations.
 2.2 Jailbreak Attacks against MLLMs
@@
 2.3 Adversarial Training for MLLMs
-Adversarial training has become one of the most effective methods to defend against escape attacks from MLLMs.
+Adversarial training [32, 33, 69, 52, 34, 25] has become one of the most effective methods to defend against escape attacks from MLLMs.
 Early training-based defense strategies mainly improve the defense capability of models through parameter fine-tuning.
@@
 However, existing adversarial training methods still face challenges in dealing with the gap between continuous representations and discrete texts.
-To address these challenges, we propose a novel end-to-end joint training framework to enhance MLLMs’ robustness against jailbreak attacks.
-Our approach first performs adversarial training on the projector by processing both original and noise-perturbed samples, optimizing through MSE minimization.
-We then innovatively integrate the projector and language model into a unified feature space, enabling robust multi-modal representations that resist adversarial perturbations.
-This joint training framework, combined with dynamic loss weighting, significantly improves the model’s defensive capabilities through cross-modal consistency and feature-level supervision.
+To address these challenges, we propose ProEAT, an end-to-end collaborative training framework designed to enhance the jailbreak attack defense capabilities of MLLMs.
+The methodology initiates with adversarial projector training, where dual-stream processing of pristine and adversarially perturbed inputs optimizes modality alignment via MSE-driven gradient descent.
+Subsequently, we establish parameter-shared coupling between the projector and language model in a unified embedding space, enabling robust multi-modal representations that resist adversarial perturbations.
+This joint training framework employs dynamic multi-task optimization that enforces cross-modal semantic consistency constraints while providing feature-level supervision of latent space geometry and attention patterns.
+Through this joint adversarial training approach with adaptive loss weighting, ProEAT achieves an average of 34% stronger defense efficacy against jailbreak attacks compared to conventional methods.
 Figure 3: Defense Framework.